Evaluate the integral of e^x + e^(-x) dx/(e^x - e^(-x))


1
Expert's answer
2021-03-30T07:35:29-0400

GIVEN:

"\int{e^x+ e^{-x}\over e^x- e^{-x}}dx"


LET,

"e^x-e^{-x}" =t

Therefore,

Differentiating this we get,


"(e^x+ e^{-x})" dx=dt


Then,


"\int{dt\over t}" We get,


"=ln(t)\\=ln(e^x-e^{-x})+c" ....answer